# Geometric QCD II: The Confining Twistor String and Meson Spectrum

### 12 Mar 2026 v3

@@ 10.8 The General Master Field $\hat{X}$ and the Word Space
However, in the planar limit ($N_{c}\rightarrow\infty$), the Master Field operates in a strictly infinite-dimensional Hilbert space. As established by the Master Field Theorem proven by the author in 1994 [30, 31], any generic parametric-invariant loop functional expandable in iterated path integrals (guaranteed to be topologically complete by Chen’s theorem [1]) can be mapped exactly onto the non-commutative moment space of Free Probability and represented exactly as a Magnus-ordered exponential path integral over a non-commutative word space.
@@
These $Q$ tensors are related to the invariant $\mathcal{W}$ tensors by a free cumulant expansion, uniquely solvable term by term for $Q(\mathcal{W})$, assuming the $\mathcal{W}$ tensors are finite. Therefore, the breakdown at the eighth order is definitively not an artifact of a restrictive algebraic ansatz. Rather, this lack of a solution for $\mathcal{W}^{(8)}$ rigorously implies that the assumption of an analytic functional $W[P]$ is violated beyond the sixth order of the Magnus expansion. The physical resolution is that the true solution must be a singular functional possessing a non-local dependence on the momentum loop $P$.
@@
Before this minimization, the area is simply a quadratic Dirichlet-Hilbert functional 3.28 governed by the singular principal-value kernel $\mathcal{H}$. However, the strict minimization over parametrizations transforms it into a fundamentally singular functional because the optimal (isothermal) parametrization becomes implicitly and non-locally tied to the geometric shape of the loop. Any local change to the shape of the loop (for instance, tracking the normalized momentum $\tilde{P}^{\prime}=P^{\prime}/\int|P^{\prime}|d\theta$) forces a global readjustment of this parametrization. This continuous global geometric shift modifies the analytic structure (though not the position) of the pole at $\theta=\theta^{\prime}$ of the Hilbert kernel on the unit circle. Such shape-dependent, non-local singularities are fundamentally incompatible with a continuous, analytic Taylor-Magnus expansion in terms of iterated path integrals.
Our exact continuum solution—the fermion determinant evaluated on the rigid Hodge-dual minimal surface—belongs precisely to this class of singular geometric functionals. The required reparametrization invariance is dynamically enforced by the null-twistor (Virasoro) constraint, which acts as the exact algebraic mechanism implementing the Douglas minimization over parametrizations. The algebraic catastrophe at $\mathcal{W}^{(8)}$ is thus the mathematical signature of the string geometry asserting its non-local rigidity and justifying the transition from a 1D algebraic Master Field to the 4D continuous geometry of the rigid twistor string.
@@ 11 The twistor parametrization of the loop space measure
As emphasized in Remark 1.2, we are not evaluating an abstract functional measure over a fluctuating 2D surface. Therefore, the measure we derive here does not encounter the ambiguities of Teichmüller space. Our starting point is the rigorously defined, linear measure of the 1D quark-loop phase space: $\int\mathcal{D}C^{\prime}\mathcal{D}P$. The transition to twistor space is strictly a local (in $\theta$) change of variables, mapping the bounding curve velocity $C^{\prime}(\theta)\propto\text{Im}(z\lambda\sigma\mu)|_{z=e^{i\theta}}$ directly to the boundary twistors $\lambda,\mu$. Because the twistors are localized at the boundary, this transformation involves no functional anomalies. Once the zero modes are fixed, we derive the corresponding twistor measure via standard, explicitly calculable algebraic matrix Jacobians.
@@ 11.1 Gauge Symmetry and the Need for Gauge Fixing
The linear integration measure $\mathcal{D}C$ (10.3) is redundant due to the reparametrization invariance of the Wilson loop $W[C]$. To define the path integral rigorously, we must fix this gauge freedom and factor the measure by the volume of diffeomorphisms of the unit circle.
@@ 11.2 The residual $U(1)$ gauge symmetry and normalized twistors
A common objection to minimal surfaces in the context of QCD strings arises from Euclidean intuition, where a surface of finite area bounded by a loop must satisfy strict embedding theorems (e.g., the Meeks-Yau theorem). In the Euclidean "rubber band" analogy, a surface cannot pass through itself without creating a physical self-intersection, which would typically be singular or unstable.
Euclidean Space: The surface is constrained to a finite sheet. Any attempt to introduce a logarithmic branching point would force the sheet to intersect itself.
Minkowski/Twistor Space: The surface behaves like a helicoid. As the trajectory winds around the twistor pole $\lambda(z)\propto 1/(z-z_{0})$ or $\mu(z)\propto 1/(z-z_{0})$, it moves to a "different sheet" in the complexified coordinate space. This is why the surface can have an infinite area in the Minkowski sense—it is effectively a Riemann surface with an infinite number of sheets: the coordinate map function $f_{\mu}(z)\propto\log(z-z_{0})$.
This multiple-valued coordinate is what allows the poles to exist without the "self-intersection" catastrophe. The surface doesn’t intersect itself for the same reason a spiral staircase doesn’t intersect the floor below it: it has moved in an orthogonal dimension. The "infinite area" is simply the total area of all these sheets, and the monodromy is the "height" gained between each floor of the staircase, which we quantize as:
@@ 12.4 Phase space path integral for quark loop amplitudes
Taking the limit of zero quark mass and focusing on the local factor dependent on $P$, we compute the integral over the local momentum variable $q_{\mu}(\theta)=P_{\mu}(\theta)d\theta$:
@@
Remarkably, the factor $e^{-3\rho}$ from the momentum integration exactly cancels the measure factor $e^{3\rho}$ derived from the twistor transformation, rendering the effective local volume element scale-invariant.
@@
The exact cancelation of the Liouville factors, $e^{3\rho}\cdot e^{-3\rho}=1$, is a necessary consequence of the symplectic geometry of the loop phase space. The canonical path integral measure $\mathcal{D}P\mathcal{D}C$ is invariant under canonical transformations, including the scaling dilatation $C_{\mu}\to sC_{\mu}$ and $P_{\mu}\to s^{-1}P_{\mu}$. In our construction, the twistor transformation introduces a specific scaling of the coordinate velocity $v\sim e^{\rho}$. The Jacobian of this transformation ($e^{3\rho}$) represents the compression of the coordinate measure. However, the integration over the conjugate momentum $P_{\mu}$ acts as an inverse scaling operator (a Fourier transform), generating the counter-factor $e^{-3\rho}$. The resulting scale invariance of the effective volume element ensures that the theory remains conformal at the classical level, consistent with the dimensionless nature of the phase space volume element $dP\wedge dC\sim\hbar$.
@@ 12.5 The Holographic Liouville Theory
In our framework, the path integral measure over loop space is rigorously constructed through a systematic procedure: we begin with a linear measure, impose the Virasoro constraint via twistor parametrization, and fix the residual local scaling symmetry. The resulting measure is manifestly real and positive. Crucially, this measure does not contain an exponential of the KKS symplectic form—a hallmark of conventional string quantization and Conformal Field Theory (CFT). We emphasize that our approach does not involve quantizing a fundamental string in the Nambu-Goto sense, but rather constitutes a faithful implementation of the solution to the Makeenko–Migdal loop equations in momentum loop space.
@@ 12.6 Block Matrix Structure and Diagonalization via Staggered Spinors
This reduction from $4\times 4$ to $2\times 2$ matrices significantly reduces the computational complexity for numerical evaluation. The trace of the remaining product of our projection operators is real (by the charge conjugation symmetry of Dirac traces), but it is not positive definite, and in general, it oscillates. It is precisely this highly oscillatory nature of the path integral that mandates its exact analytical evaluation via Picard-Lefschetz theory and Catastrophe Theory.
@@ 13.1 Twistor Holography (Gauge Holography)
Because the path integral over the twistor target space is highly oscillatory, one might assume that extracting the physical spectrum requires massive stochastic lattice simulations. However, by analyzing the holomorphic structure of the effective action, we find that the discrete mass spectrum can be extracted exactly and analytically. Planar QCD ceases to be a theory of stochastic quantum fluctuations; it is a theory of pure classical complex geometry.
@@ 14.2.1 Spinor Fields
Comparing this to our spectral expansion, the Fourier coefficients $L_{n}$ are exactly determined as algebraic superpositions of the pole parameters:
@@
Consequently, transforming the functional measure from the $N$ discrete boundary values to the physical moduli $(R_{i},z_{i})$ involves exactly two finite, calculable matrix Jacobians:1. The constant Jacobian of the Discrete Fourier Transform matrix (the Fourier matrix $e^{2\pi\mathrm{i}kn/N}$), which maps the boundary values to the Fourier coefficients $L_{n}$.2. The purely algebraic Jacobian mapping the Fourier coefficients to the pole parameters $(R_{i},z_{i})$.We do not explicitly compute these algebraic Jacobians here, because they contribute strictly to the pre-exponential fluctuation determinant of the path integral. For the extraction of the exact geometric mass spectrum, the relevant physics is governed entirely by the stationary phase (the saddle-point equations) and the topological monodromies of the effective action. As we will demonstrate, the S-matrix poles emerge from the constructive interference of the geometric series over the topological winding number $k$. In this resonant limit, the multi-valued action scales as $k\Delta S_{eff}$. Consequently, the transverse quantum fluctuations around the saddle point are dynamically damped by a factor of $1/k$. This perfectly crushes the pre-exponential factor, rendering the explicit form of the Jacobian irrelevant to the exact localization of the S-matrix poles. However, the explicit existence of this algebraic mapping mathematically guarantees that our twistor measure is strictly finite and constructible, firmly rooting the evaluation of the path integral in standard finite-dimensional complex analysis rather than abstract string-theoretic functional geometry.
@@ 14.2.2 Liouville Field
The Liouville field $\rho(z,\bar{z})$ is extended inside the unit circle via an exact formula
@@ 14.3 The Boundary Terms in the action
With the restored measure factor, the power is $p=n+m+1$. Since $n,m\geq 0$, the denominator $n+m+1\geq 1$ is strictly positive, ensuring there are no singularities in the complex plane.
@@ 14.4 Stochastic Quantization and Complex Langevin Dynamics
The exact mathematical evaluation of this highly oscillatory complexified path integral is rigorously achieved via Stochastic Quantization. The quantum statistical evolution of the twistor string is governed by the Complex Langevin Equation (CLE) with holomorphic drift forces derived from the total spectral action $S_{total}=S_{q}+S_{\gamma}+S_{T}+S_{\rho}+S_{\mathcal{O}}$. We emphasize that the CLE is utilized here not as a numerical simulation technique, but as an exact continuous mathematical reformulation of the quantum path integral.
@@
    Crucially, the noise terms $\nu_{\xi}$ and $\nu_{\tilde{\xi}}$ are independent complex Gaussian variables representing the exact quantum fluctuations of the twistor fields.
@@
    Constraint Evolution ($\gamma_{k}$): The multipliers evolve to enforce the exact algebraic constraints:
@@
In this mathematical framework, the deterministic drift terms ($\nabla S$) represent the exact holomorphic gradient flow in complexified phase space. In Picard-Lefschetz theory, the union of all these deterministic flow lines originating from a complex saddle point constructs a middle-dimensional real manifold known as a Lefschetz thimble. If the noise $\nu$ were zero, the system would trivially collapse onto a 1D classical trajectory (the WKB approximation). However, the multidimensional stochastic noise $\nu$ dynamically forces the probability distribution to completely explore the transverse width of these $N$-dimensional thimbles, strictly incorporating all exact quantum fluctuations (the “vicinity” of the saddle points).
The phenomenon of exact S-matrix quantization we are looking for corresponds to a topological bifurcation of this stochastic system. At certain critical resonance energies, the complex Hessian develops a null space and the classical restoring drift strictly vanishes along a continuous topological valley. Unimpeded by any restoring gradient force, the stochastic noise $\nu$ drives the complex trajectory to infinity along this flat zero mode. This infinite flat-valley integration geometrically generates the exact macroscopic S-matrix pole of the path integral, governed by Catastrophe Theory. Before studying this phenomenon in general form, we consider the simple solvable limit: the WKB quantization corresponding to the large rotational momentum of our rigid string.
@@ 15.1 Twisted Boundary Conditions, Symmetrization, and Spin Projections
Before solving the exact spectral equations, we justify the choice of the twisted geometry from the perspective of the semiclassical quark path integral. In Planar QCD, the effective Minkowski action involves a sum over quark trajectories weighted by the area of the confining minimal surface.
To extract the spectrum of states with macroscopic angular momentum $J$, we impose twisted boundary conditions. We require the quark positions at the ends of the string to undergo a physical spatial rotation by an angle $\phi$ over the period of the motion, with the angular momentum $J$ acting as the thermodynamic conjugate to this twist. The minimization of the boundary length for a rotating particle yields a helical trajectory in Minkowski spacetime. The minimal surface spanning two twisted helical boundaries is the helicoid, as illustrated in Figure 14.
@@
In our twistor string formulation, we project this twisted kinematics directly into the boundary twistor data. This is achieved by introducing the $SU(2)_{L}$ and $SU(2)_{R}$ projections, integrating over the target space rotation angle $\phi$ with the canonical spin weight:
$\mathcal{P}_{J}=\oint d\phi\exp\left(\mathrm{i}\phi J\right)$ (15.1)
When the macroscopic boundary of the string rotates by an angle $\phi$, the boundary twistors must geometrically twist. To enforce this, we introduce a fractional winding parameter $a$ corresponding to twisted boundary conditions for $\lambda(z)$.
Furthermore, because macroscopic QCD rigorously conserves parity, the physical effective action must be symmetrized over the Self-Dual (SD) and Anti-Self-Dual (ASD) minimal surfaces. In the twistor parameterization, this corresponds to activating the twisted boundary condition for $\lambda(z)$ while keeping the mirror sector $\mu(z)$ frozen to a constant basis spinor, and adding a mirror term with $\mu(z)$ to the exponent. The Hodge-duality (or anti-duality) explicitly breaks the 3D embedding constraint, forcing the chiral surface to genuinely span 4-dimensional Minkowski space.
To analytically generate the perfect resonance leading to the exact WKB saddle point, we must consider the limit of an infinitely large number of geometric windings $k\rightarrow\infty$. As the integration variable traverses the flat valley $k$ times, the accumulated action scales strictly linearly with $k$. Crucially, this means the transverse Hessian scales as $\mathcal{H}_{k}\propto k$. In the asymptotic resonance limit $k\rightarrow\infty$, the transverse restoring force becomes strictly infinite. The stochastic variance of the quantum fluctuations scales as $1/k\rightarrow 0$. The path integral dynamically localizes to the exact WKB saddle point, completely shielding the topological monodromy from quantum noise. These heuristic arguments will be made explicit and quantitative in this section, which will end with exact WKB quantization in a topological sector with only one pole of twistor field.

### 23 Mar 2026 v4

@@ 10.8 The General Master Field $\hat{X}$ and the Word Space
However, in the planar limit ($N_{c}\rightarrow\infty$), the Master Field operates in a infinite-dimensional Hilbert space. As established by the Master Field Theorem proven by the author in 1994 [26, 27], any generic parametric-invariant loop functional expandable in iterated path integrals (guaranteed to be topologically complete by Chen’s theorem [1]) can be mapped onto the non-commutative moment space of Free Probability and represented as a Magnus-ordered exponential path integral over a non-commutative word space.
@@
These $Q$ tensors are related to the invariant $\mathcal{W}$ tensors by a free cumulant expansion, uniquely solvable term by term for $Q(\mathcal{W})$, assuming the $\mathcal{W}$ tensors are finite. Therefore, the breakdown at the eighth order is definitively not an artifact of a restrictive algebraic ansatz. Rather, this lack of a solution for $\mathcal{W}^{(8)}$ implies that the assumption of an analytic functional $W[P]$ is violated beyond the sixth order of the Magnus expansion. The physical resolution is that the true solution must be a singular functional possessing a non-local dependence on the momentum loop $P$.
@@
Before this minimization, the area is simply a quadratic Dirichlet-Hilbert functional 3.22 governed by the singular principal-value kernel $\mathcal{H}$. However, the strict minimization over parametrizations transforms it into a fundamentally singular functional because the optimal (isothermal) parametrization becomes implicitly and non-locally tied to the geometric shape of the loop. Any local change to the shape of the loop (for instance, tracking the normalized momentum $\tilde{P}^{\prime}=P^{\prime}/\int|P^{\prime}|d\theta$) forces a global readjustment of this parametrization. This continuous global geometric shift modifies the analytic structure (though not the position) of the pole at $\theta=\theta^{\prime}$ of the Hilbert kernel on the unit circle. Such shape-dependent, non-local singularities are fundamentally incompatible with a continuous, analytic Taylor-Magnus expansion in terms of iterated path integrals.
Our continuum solution—the fermion determinant evaluated on the rigid Hodge-dual minimal surface—belongs to this class of singular geometric functionals. The required reparametrization invariance is dynamically enforced by the null-twistor (Virasoro) constraint, which acts as the algebraic mechanism implementing the Douglas minimization over parametrizations. The algebraic inconsistency at $\mathcal{W}^{(8)}$ is thus the mathematical signature of the string geometry asserting its non-local rigidity and justifying the transition from a 1D algebraic Master Field to the 4D continuous geometry of the rigid twistor string.
@@ 11 The twistor parametrization of the loop space measure
As emphasized in Remark 1.2, we are not evaluating an abstract functional measure over a fluctuating 2D surface. Therefore, the measure we derive here does not encounter the ambiguities of Teichmüller space. Our starting point is the well defined, linear measure of the 1D quark-loop phase space: $\int\mathcal{D}C^{\prime}\mathcal{D}P$. The transition to twistor space is a local (in $\theta$) change of variables, mapping the bounding curve velocity $C^{\prime}(\theta)\propto\text{Im}(z\lambda\sigma\mu)|_{z=e^{i\theta}}$ directly to the boundary twistors $\lambda,\mu$. Because the twistors are localized at the boundary, this transformation involves no functional anomalies. Once the zero modes are fixed, we derive the corresponding twistor measure via standard, explicitly calculable algebraic matrix Jacobians.
@@ 11.1 Gauge Symmetry and the Need for Gauge Fixing
The linear integration measure $\mathcal{D}C$ (10.3) is redundant due to the reparametrization invariance of the Wilson loop $W[C]$. To define the path integral without ambiguity, we must fix this gauge freedom and factor the measure by the volume of diffeomorphisms of the unit circle.
@@ 11.2 The residual $U(1)$ gauge symmetry and normalized twistors
In Euclidean signature, minimal surfaces of finite area bounded by a loop satisfy embedding theorems (e.g., the Meeks-Yau theorem) that preclude self-intersections. In Minkowski twistor space, however, the surface behaves like a helicoid. As the trajectory winds around the twistor pole $\lambda(z)\propto 1/(z-z_{0})$ or $\mu(z)\propto 1/(z-z_{0})$, it moves to a "different sheet" in the complexified coordinate space. This is why the surface can have an infinite area in the Minkowski sense—it is effectively a Riemann surface with an infinite number of sheets: the coordinate map function $f_{\mu}(z)\propto\log(z-z_{0})$.
This multiple-valued coordinate is what allows the poles to exist without the "self-intersection" inconsistency. The surface doesn’t intersect itself for the same reason a spiral staircase doesn’t intersect the floor below it: it has moved in an orthogonal dimension. The "infinite area" is simply the total area of all these sheets, and the monodromy is the "height" gained between each floor of the staircase, which we quantize as:
@@ 12.4 Phase space path integral for quark loop amplitudes
Taking the limit of zero quark mass (see the remark 16.2 for the definition of our effective quark mass) and focusing on the local factor dependent on $P$, we compute the integral over the local momentum variable $q_{\mu}(\theta)=P_{\mu}(\theta)d\theta$:
@@
Remarkably, the factor $e^{-3\rho}$ from the momentum integration precisely cancels the measure factor $e^{3\rho}$ derived from the twistor transformation, rendering the effective local volume element scale-invariant.
@@
The cancelation of the Liouville factors, $e^{3\rho}\cdot e^{-3\rho}=1$, is a necessary consequence of the symplectic geometry of the loop phase space. The canonical path integral measure $\mathcal{D}P\mathcal{D}C$ is invariant under canonical transformations, including the scaling dilatation $C_{\mu}\to sC_{\mu}$ and $P_{\mu}\to s^{-1}P_{\mu}$. In our construction, the twistor transformation introduces a specific scaling of the coordinate velocity $v\sim e^{\rho}$. The Jacobian of this transformation ($e^{3\rho}$) represents the compression of the coordinate measure. However, the integration over the conjugate momentum $P_{\mu}$ acts as an inverse scaling operator (a Fourier transform), generating the counter-factor $e^{-3\rho}$. The resulting scale invariance of the effective volume element ensures that the theory remains conformal at the classical level, consistent with the dimensionless nature of the phase space volume element $dP\wedge dC\sim\hbar$.
@@ 12.5 The Holographic Liouville Theory
In our framework, the path integral measure over loop space is constructed through a systematic procedure: we begin with a linear measure, impose the Virasoro constraint via twistor parametrization, and fix the residual local scaling symmetry. The resulting measure is manifestly real and positive. Crucially, this measure does not contain an exponential of the KKS symplectic form—a hallmark of conventional string quantization and Conformal Field Theory (CFT). We emphasize that our approach does not involve quantizing a fundamental string in the Nambu-Goto sense, but rather constitutes a faithful implementation of the solution to the MM loop equations in momentum loop space.
@@ 12.6 Block Matrix Structure and Diagonalization via Staggered Spinors
This reduction from $4\times 4$ to $2\times 2$ matrices significantly reduces the computational complexity for numerical evaluation. The trace of the remaining product of our projection operators is real (by the charge conjugation symmetry of Dirac traces), but it is not positive definite, and in general, it oscillates. It is this oscillatory nature of the path integral that mandates its analytical evaluation via Picard-Lefschetz theory and Catastrophe Theory.
@@ 13.1 Twistor Holography (Gauge Holography)
Because the path integral over the twistor target space is highly oscillatory, one might assume that extracting the physical spectrum requires massive stochastic lattice simulations. However, by analyzing the holomorphic structure of the effective action, we find that the discrete mass spectrum can be extracted analytically. Our solution to Planar QCD is not a theory of quantum fluctuations in any space-time dimensions; it is a problem of classical geometry.
@@ 14.2.1 Spinor Fields
Comparing this to our spectral expansion, the Fourier coefficients $L_{n}$ are determined as algebraic superpositions of the pole parameters:
@@
Consequently, transforming the functional measure from the $N$ discrete boundary values to the physical moduli $(R_{i},z_{i})$ involves two finite, calculable matrix Jacobians:1. The constant Jacobian of the Discrete Fourier Transform matrix (the Fourier matrix $e^{2\pi\mathrm{i}kn/N}$), which maps the boundary values to the Fourier coefficients $L_{n}$.2. The purely algebraic Jacobian mapping the Fourier coefficients to the pole parameters $(R_{i},z_{i})$.We do not explicitly compute these algebraic Jacobians here, because they contribute to the pre-exponential fluctuation determinant of the path integral. For the extraction of the geometric mass spectrum, the relevant physics is governed entirely by the stationary phase (the saddle-point equations) and the topological monodromies of the effective action. As we will demonstrate, the S-matrix poles emerge from the constructive interference of the geometric series over the topological winding number $k$. In this resonant limit, the multi-valued action scales as $k\Delta S_{eff}$. Consequently, the transverse quantum fluctuations around the saddle point are dynamically damped by a factor of $1/k$. This exponential part dominates over the pre-exponential factor, rendering the explicit form of the Jacobian irrelevant to the exact localization of the S-matrix poles. However, the explicit existence of this algebraic mapping mathematically guaranties that our twistor measure is finite and constructible, firmly rooting the evaluation of the path integral in standard finite-dimensional complex analysis rather than abstract string-theoretic functional geometry.
@@ 14.2.2 Liouville Field
The Liouville field $\rho(z,\bar{z})$ is extended inside the unit circle via an formula
@@ 14.3 The Boundary Terms in the action
With the restored measure factor, the power is $p=n+m+1$. Since $n,m\geq 0$, the denominator $n+m+1\geq 1$ is positive, ensuring there are no singularities in the complex plane.
@@ 14.4 Stochastic Quantization and Complex Langevin Dynamics
The mathematical evaluation of this highly oscillatory complexified path integral is achieved via Stochastic Quantization. The quantum statistical evolution of the twistor string is governed by the Complex Langevin Equation (CLE) with holomorphic drift forces derived from the total spectral action $S_{total}=S_{q}+S_{\gamma}+S_{T}+S_{\rho}+S_{\mathcal{O}}$. We emphasize that the CLE is utilized here not as a numerical simulation technique, but as an continuous mathematical reformulation of the quantum path integral.
@@
    Crucially, the noise terms $\nu_{\xi}$ and $\nu_{\tilde{\xi}}$ are independent complex Gaussian variables representing the quantum fluctuations of the twistor fields.
@@
    Constraint Evolution ($\gamma_{k}$): The multipliers evolve to enforce the algebraic constraints:
@@
In this mathematical framework, the deterministic drift terms ($\nabla S$) represent the holomorphic gradient flow in complexified phase space. In Picard-Lefschetz theory, the union of all these deterministic flow lines originating from a complex saddle point constructs a middle-dimensional real manifold known as a Lefschetz thimble. If the noise $\nu$ were zero, the system would trivially collapse onto a 1D classical trajectory (the WKB approximation). However, the multidimensional stochastic noise $\nu$ dynamically forces the probability distribution to completely explore the transverse width of these $N$-dimensional thimbles, incorporating all quantum fluctuations (the “vicinity” of the saddle points).
The phenomenon of exact S-matrix quantization we are looking for corresponds to a topological bifurcation of this stochastic system. At certain critical resonance energies, the complex Hessian develops a null space and the classical restoring drift vanishes along a continuous topological valley. Unimpeded by any restoring gradient force, the stochastic noise $\nu$ drives the complex trajectory to infinity along this flat zero mode. This infinite flat-valley integration geometrically generates the exact macroscopic S-matrix pole of the path integral, governed by Catastrophe Theory. Before studying this phenomenon in general form, we consider the simple solvable limit: the WKB quantization corresponding to the large rotational momentum of our rigid string.
@@ 15.1 Twisted Boundary Conditions, Symmetrization, and Spin Projections
Before solving the spectral equations, we justify the choice of the twisted geometry from the perspective of the semiclassical quark path integral. In Planar QCD, the effective Minkowski action involves a sum over quark trajectories weighted by the area of the confining minimal surface.
To extract the spectrum of states with macroscopic angular momentum $J$, we impose twisted boundary conditions. We require the whole string to undergo a physical spatial rotation by an angle $\alpha$ over the period of the motion, with the angular momentum $J$ acting as the thermodynamic conjugate to this twist. The minimization of the boundary length for a rotating particle yields a helical trajectory in Minkowski spacetime. The minimal surface spanning two twisted helical boundaries is the helicoid, as illustrated in Figure 14.
@@
In our twistor string formulation, we project this twisted kinematics directly into the boundary twistor data. This is achieved by introducing the $SU(2)_{L}$ and $SU(2)_{R}$ projections, integrating over the target space rotation angle $\alpha$ with the canonical spin weight:
$\hat{P}_{J}A=\sum_{k}\oint d\alpha\exp\left(-\mathrm{i}k\alpha J\right)A_{k}(\alpha)$ (15.1)
Here $A_{k}(\alpha)$ is the amplitude after $k$ winding around the quark loop, with the boundary condition that the rotation angle $\alpha$ is introduced at every full circle around the quark loop. When the macroscopic boundary of the string rotates by an angle $\alpha$, the boundary twistors must geometrically twist. To enforce this, we introduce a fractional winding parameter $a$ corresponding to twisted boundary conditions for $\lambda(z),\mu(z)$.
To analytically generate the perfect resonance leading to the exact WKB saddle point, we must consider the limit of an infinitely large number of geometric windings $k\rightarrow\infty$. As the integration variable traverses the flat valley $k$ times, the accumulated action scales linearly with $k$. Crucially, this means the transverse Hessian scales as $\mathcal{H}_{k}\propto k$. In the asymptotic resonance limit $k\rightarrow\infty$, the transverse restoring force becomes infinite. The stochastic variance of the quantum fluctuations scales as $1/k\rightarrow 0$. The path integral dynamically localizes to the exact WKB saddle point, completely shielding the topological monodromy from quantum noise. These heuristic arguments will be made explicit and quantitative in this section, which will end with exact WKB quantization in a topological sector with only one pole of the twistor field.